# An alternative theory of magnetic flux tubes in strong fields via axion origin photons

### v1

@@ 1 Introduction
On the other hand, we understand that the existence of the Parker-Biermann cooling effect is associated with the so-called thermomagnetic Ettingshausen-Nernst effect (see Apendix A in V. D. Rusov et al. (2021)). Due to the large temperature gradient in the tachocline, the thermomagnetic EN effect (A. V. Ettingshausen & W. Nernst, 1886; E. H. Sondheimer, 1948; L. J. Spitzer, 1956; Y. B. Kim & M. J. Stephen, 1969) creates electric currents that are inversely proportional to the strong magnetic field of the tachocline. As we showed earlier (V. D. Rusov et al., 2021), the toroidal magnetic field of tachocline by means of the thermomagnetic Ettingshausen-Nernst effect (Appendix A in V. D. Rusov et al. (2021)) “neutralizes” the magnetic field of the solar core $\sim 5\cdot 10^{7}~G$ (W. A. Fowler et al., 1955; S. Couvidat et al., 2003). It means that, using the thermomagnetic EN effect, a simple estimate of the magnetic pressure of an ideal gas in the tachocline of e.g. the Sun,
@@
can indirectly prove that by using the holographic principle of quantum gravity (see Apendix C in V. D. Rusov et al. (2021)), the repelling toroidal magnetic field of the tachocline exactly “neutralizes” the magnetic field in the Sun’s core (see Fig. 2, and also Fig. A1 in V. D. Rusov et al. (2021))
$B_{tacho}^{Sun}=4.1\cdot 10^{7}~G=-B_{core}^{Sun},$ (2)
where the projections of the magnetic fields of the tachocline and the core have equal values but opposite directions.
Surprisingly, our experimental data provide observational evidence in support of a two-dimensional holographic explanation of the Sun and a powerful key to understanding the physics of quantum gravity in the Universe! This is explained once again by the fact that the experimental data, in which the toroidal magnetic field of the tachocline exactly ”neutralizes” the magnetic field in the solar core, should also theoretically describe both a two-dimensional holographic explanation of the solar tachocline, in which it is the magnetic field that must be solved, either in conformal field theory (see e.g. analogs in (R. Boyack et al., 2024; T. Kimura & S. Ozaki, 2019; A. B. Zamolodchikov, 1989; C. Callan et al., 1995; L. Hu et al., 2024; C. Cartwright et al., 2022; H. Nastase & K. Skenderis, 2020; V. A. Miransky & I. A. Shovkovy, 2015)) or at least in quantum mechanics (see e.g. analogs in (U. Gürsoy, 2021; S. Heefer, 2024)), and a three-dimensional magnetic field via gravity, which, in contrast to pseudo-Riemannian geometry, applies the extended (beyond Einstein geometry) Finsler geometry (see (U. Gürsoy, 2021; S. Heefer, 2024)), for example, inside the solar tachocline.
In other words, the “encoded” two-dimensional information about the magnetic field of the tachocline in conformal field theory corresponds to three-dimensional information about the magnetic field in Finsler gravity, which is reduced to the curvature of space-time in a three-dimensional volume with the radius of the tachocline of the Sun.
Such strong magnetic fields in a tachocline of e.g. the Earth, Sun, magnetic white dwarfs, accreting neutron stars and black holes can predict the exact “neutralization” of the magnetic field in the core of these stars and in a black hole (see Appendix C in V. D. Rusov et al. (2021)).
From here we understand that when space-time curvature appears, gravitational waves arise, which ultimately uncover, as not surprisingly, the more powerful Finslerian nature of quantum gravity. So Finslerian quantum gravity can be applied in our case, as we know, to holograms of the Universe, black holes, stars and planets, for example, the Sun.
Based on the “experimental example” of the existence and understanding of the physics of the holographic principle of quantum gravity, we apply the famous result of Stephen Hawking, who at a conference in Stockholm on August 28, 2015 announced that he had solved the information paradox. According to Hawking (S. W. Hawking, 2015a, b; S. W. Hawking et al., 2016), the AdS/CFT correspondence (see (J. Maldacena, 1998)) showed no loss of information, where all information about matter and energy in the three-dimensional volume of the black hole is encoded in the form of a hologram on its two-dimensional surface of the event horizon (see (G. ’t Hooft, 1993; L. Susskind, 1995; J. Maldacena, 2005; M. Hanada et al., 2014)). It resembles an analogue of the famous photohologram (see (D. Gabor, 1948)), which is a certain type of photograph that produces a three-dimensional image under good lighting; all the information describing a three-dimensional scene is encoded in a pattern of light and dark areas applied to a two-dimensional film. This means that the possible description and solution of the ”encoded” holographic two-dimensional ”film” is predetermined by the fact that the problem of the information paradox, known as the apparent contradiction between general relativity and quantum mechanics, completely disappears. This approach ultimately leads to a holographic principle of quantum gravity, which can be generalized to any physical system occupying space-time. For example, in the holograms of the Universe (N. Afshordi et al., 2017), black holes, white dwarfs, neutron stars or, surprisingly, in the hologram of the Sun’s tachocline!
Let us note one more important mechanism of quantum gravity (see Fig. 2b). The most surprising fact is that the antidynamo model of a magnetic flux tube at strong fields of the order of $10^{7}~G$ is a consequence of both the thermomagnetic EN effect and the so-called solar holographic antidynamo, in which the poloidal field arises directly from the toroidal field, but not vice versa (see Fig. 2b; and also Fig. C.1a in V. D. Rusov et al. (2021)).
Using the quantum gravity mechanism, and hence, the strong MFT field, we are interested in the existence of dark matter axions identical to solar axions, which strongly affects the the conversion of axion origin photons in magnetic O-loops near the tachocline, and are thus related to the so-called thermomagnetic Ettingshausen-Nernst effect through the Parker-Biermann cooling effect (see Apendix A in V. D. Rusov et al. (2021)).
@@
Figure 1: An illustration of the main possible processes of a magnetically active star of the Sun type. (a) $\alpha$-effect, $\Omega$-effect and BL mechanism as components of the solar dynamo model. The $\Omega$-effect (blue) depicts the transformation of the primary poloidal field into a toroidal field by differential rotation. Regeneration of the poloidal field is then performed either by the $\alpha$-effect (top) or by the BL mechanism (yellow in the middle). In case of $\alpha$-effect, the toroidal field at the base of the convection zone is subject to cyclonic turbulence. In the BL mechanism, the main process of regeneration of the poloidal field (based on the $\Omega$-effect (blue)) is the formation of sunspots on the surface of the Sun from the rise of buoyant toroidal flux tubes from the base of the convection zone. The magnetic fields of these sunspots closest to the equator in each hemisphere diffuse and join, and the field due to the spots closer to the poles has a polarity opposite to the current that initiates rotation of the polarity. The newly formed polar magnetic flux is transported by the meridional flow to deeper layers of the convection zone, thereby creating a new large-scale poloidal field. Derived from S. Sanchez et al. (2014). (b) 11-year variations of sunspot magnetic fields as a component of the solar anti-dynamo. In contrast to the component of the solar dynamo model (a) [^2] If the magnetic field at these levels is maintained by a dynamo, and the magnetic fields are diffuse, then the field strength is unlikely to be more than $\sim 10^{5}~G$, because larger fields would severely inhibit dynamo action (see E. E. Deluca & P. A. Gilman (1986))., the mechanism of the fundamental holographic principle of quantum gravity (see H. Nastase & K. Skenderis (2020); N. Afshordi et al. (2017)), in which the hologram of the Sun is encoded in the magnetic field on its two-dimensional surface tachocline and is equivalent to the magnetic field (from the core to the tachocline) in the internal three-dimensional space of the Sun’s gravity, and consequently, the formation of the thermomagnetic EN effect (see L. J. Spitzer (1962, 2006); V. Rusov et al. (2015)), emphasizes that this process is associated with the continuous transformation of toroidal field into poloidal one ($T\rightarrow P$ transformation), but not vice versa ($P\rightarrow T$). For this purpose, we note (seeV. D. Rusov et al. (2021)) that the modulations of the ADM halo in the Galactic Center (GC), which is an indicator of the periods of S-stars, e.g. S-102 with a period of about 11 years, lead (via vertical density waves from the black hole to the solar neighborhood) to modulations of the ADM density, gravitationally trapped in the interior of the Sun. This means that anticorrelation modulations of the ADM density control the solar luminosity or sunspot cycles, in which the sunspot variations are identical to the 11-year variations of the magnetic flux tubes (see Section 2): maximum (red lines) and minimum (green-red lines).
@@
Taking into account the above, we structure our paper as follows. In Section 2, we discuss the physics of nearly empty flux tubes and the connection with dark matter axions out of the solar core. In Section 2.1, we present convective heating and magnetic buoyancy of flux tubes by means of axion origin photons. In Section 2.2, we discuss the physics of primary and secondary reconnection of flux tubes in the lower layers and the features of the Joy’s law tilt angle. In Section 2.2.2, we discuss the physics of the observed double maxima of sunspot cycles, where the first maximum does not depend on the tilt angle of the Joy’s law, while the second one definitely depends on the tilt angle of the Joy’s law. We also showed the relationship between the ADM density in the solar interior (or, for example, the number of anticorrelated sunspots) and the Gnevyshev gap. Finally, in Section 3, we provide a summary and outlook for this paper.
2 Physics of the theory of practically empty magnetic flux tubes at strong fields, $\sim 10^{7}~G$, at which there is a free path for photons of axion origin from the tachocline to the photosphere.
@@
Conversely, if the question arises whether there is a satisfactory alternative theory against the dynamo action, our short answer is ”yes”. We understand that the strong toroidal magnetic field in the tachocline is generated not by the dynamo action, but by the holographic mechanism of the tachocline (see Fig. 2b), which is a consequence of the holographic principle of quantum gravity in the Universe, in the black hole of our galaxy, and therefore on the Sun.
In order for our alternative theory (for example, magnetic flux tubes in strong fields using photons of axion origin) to solve all its problems (see Sections 2.1-2.2), it is necessary to solve the most important problem: In what way come sunspots, generated by a strong magnetic field in a flux tube from the base of the convection zone to the surface of the Sun, have 11-year variations without any dynamo?
The answer is very simple! Almost all physicists, and in particular dynamo physicists, know that in the Universe, and, of course, in our galaxy, there are dark matters, and the highest density of dark matter is located exactly at the center of the Galaxy, i.e. near the black hole. At the same time, we know that between the black hole and the disk there are also so-called amazing S-stars [^3] Let us remind from where the fairly young S-stars (Fig. 8a in (V. D. Rusov et al., 2021)) appear in our galaxy between the black hole and the disk. Since we know that in our galaxy, in addition to the supermassive black hole (SMBH), there is the so-called intermediate black hole (IMBH; see evidence (S. Takekawa et al., 2019)). If we take into account the evolution of binary SMBH-IMBH, then when one body (SMBH) exchanges partners of binary stars orbiting the IMBH (Fig. 8b in V. D. Rusov et al. (2021)), and through the extreme gravitational-tidal field, one star is captured (SMBH) and loses energy, while the other escapes, receives all this energy and, due to the hypervelocity of B-type stars – the IMBH sling (Fig. 8b in V. D. Rusov et al. (2021)), simply flies out of the galaxy. From here it becomes clear how in other galaxies a young S-star can end up between a black hole and a disk., which have a little more than 30 stars. The most interesting thing is that the closest S-star to the black hole is called S-102, which has an 11-yer period of around the black hole (see Fig. 6 and Fig. 8 in V. D. Rusov et al. (2021)). From here, we understand that when the period of S-102 is very close to the black hole, then the high speed of this star is such that the dark matter is practically not captured by this star. But when the elliptical star S-102 rotates farther from the black hole, then the star gravitationally captures these dark matters.
@@
Another intriguing question arises here. Unlike our galaxy, such S-type stars (between the black hole and the disk) may simply not exist in many galaxies, and consequently, there may not be any solar dynamo at all – for example, there is no Sun-like dynamo on the active star $\zeta$ Andromedae, which is part of the Zeta Andromedae star system in the constellation Andromeda (R. M. Roettenbacher et al., 2016). (It is located approximately 189 light years from Earth.) When various stars and clusters (e.g. Galactic Globular Clusters) have strong toroidal magnetic fields of $\geqslant 10^{7}~G$ in the corresponding tachoclines, then, unlike the self-excited dynamo, the manifestation of the tachocline itself in various stars and clusters is a result of the funda-mental holographic principle of quantum gravity (see J. M. Maldacena (1999); J. Maldacena (2005); S. W. Hawking (2015a, b); S. W. Hawking et al. (2016); and also N. Afshordi et al. (2017) and Fig.A.1 and Fig.C.1 in V. D. Rusov et al. (2021)). This means that the self-excited dynamo in the tachocline may not exist at all!
@@
Figure 3: Topological effects of magnetic reconnection inside the magnetic tubes with the “magnetic steps” (Fig. B.1 in V. D. Rusov et al. (2021)). The left panel shows the temperature and pressure change along the radius of the Sun from the tachocline to the photosphere (J. Bahcall & M. Pinsonneault, 1992), $L_{MS}$ is the height of the magnetic shear steps. At $R\sim 0.72~R_{Sun}$ the vertical magnetic field, which is developed from the horizontal part of the magnetic field (step) with the participation of the O-loop through the well-known Kolmogorov turbulent cascade (see Fig. 4), reaches $B_{z}\sim 3600$ T, and the magnetic pressure $p_{ext}=B^{2}/8\pi\simeq 5.21\cdot 10^{13}~erg/cm^{3}$ (J. Bahcall & M. Pinsonneault, 1992). The very cool regions along the entire convective zone caused by the Parker-Biermann cooling effect have the virtually zero internal gas pressure, i.e. the maximum magnetic pressure in the magnetic tubes. The narrow “purple” rings between the O-loop and the tube walls ($\rho_{ext}=\rho_{int}^{0}\gg\rho_{int}$) with the Parker-Bierman cooling effect inside, are a very important result of the existence of convective heating $(dQ/dt)_{2}$ in Section 2.1.
On the other hand, high-energy photon fluxes coming from the radiation zone through the thin “ring” part of the magnetic tube, i.e. between the tachocline and the “ring” (or more precisely between the magnetic wall and the O-loop of the tube; see Figs. 3 and 4), and which is applied without the Parker-Biermann cooling effect, allows to determine the convective heating $(dQ/dt)_{2}$ (see Sect. 2.1).
@@
One of the final goals of this section is to determine the general regularities of the theory of magnetic flux tubes in strong fields of the order of $\sim 10^{7}~G$ using photons of axion origin – the so-called universal Ballegooijen-Fan-Fisher model (Sec. 2.1), which are generated by the magnetic buoyancy of practically empty tubes, and as a consequence, their rise from the tachocline to the solar surface (Figs. 3-5). Another goal is the essence of the physics of the primary and secondary reconnection of magnetic flux tubes in the lower layers of the tachocline (Sect. 2.2), which is directly related to the observed features of the cycle of both double sunspot maxima (see Sect. 2.2.2) and the tilt angle of Joy’s law (see Sect. 2.2.1). Moreover, both effects (see Sect. 2.1-2.2) are caused by the existence (of dark matter?) – solar axions generated in the core of the Sun.
2.1 Radiative $(dQ/dt)_{1}$ sec-reconnectionand convective $(dQ/dt)_{2}$ heatings and the buoyant rise of magnetic flux tubes and the phenomenon of solar axions
@@
at which we have a very important (see section 2.2.1) value of
@@
at which we have a very important (see Sections 2.1-2.2) value of
@@
Using simple calculations of equations (3) and (2.1) for MFTs, it is easy to show that with the help of the
@@
Ultimately, we call these Eqs. (2.1)-(18), in honor of these remarkable “solar” physicists, the universal van Ballegooijen-Fan-Fisher model (vanBFF model).
@@
where we used the MFT heating rate which consists of the radiative heating rate $(dQ/dt)_{1}$ and convective heating rate $(dQ/dt)_{2}$ (see Eqs. (2.1)–(19), Eqs. (13) and (19) in (Y. Fan & G. Fisher, 1996) and Eqs. (8)–(9) in (M. A. Weber & Y. Fan, 2015)); the approximate ratio $|\delta T|/T_{ext}\sim 1/\beta$; the parameter $\alpha_{1}^{2}\approx 5.76$ (Y. Fan & G. Fisher, 1996; M. A. Weber & Y. Fan, 2015); $F_{rad}$ is the radiative energy flux (see Eq. (15) in (Y. Fan & G. Fisher, 1996), and also Eq. (6) and Fig. 1 in (M. A. Weber & M. K. Browning, 2016)); $\nabla_{e}\sim 1.287\nabla_{rad}$ (see Table 2 in (H. C. Spruit, 1977)); $H_{p}/a$ is the factor for the lower convection zone (see (R. M. Roettenbacher et al., 2016)), where it was previously believed that $a=(\Phi/\pi B_{tacho})^{1/2}\leqslant 0.1H_{p}$ (see 3.2.1 in (R. M. Roettenbacher et al., 2016)) with an average value of typical magnetic flux of $\Phi\sim 10^{20}-10^{22}~Mx$ (Y. Fan et al., 1993).
@@
Recalling the words of Yuhong Fan that “it remains to be more thoroughly investigated” to what extent is strong toroidal magnetic field, control by means of the self-generating dynamo in the overshoot tachocline region, one must also remember the existence of the amazing 11-year variation of dark matter density (see e.g. V. D. Rusov et al. (2021)), and, for example, the holographic mechanism of quantum gravity (see S. W. Hawking (2015b, a); S. W. Hawking et al. (2016); V. D. Rusov et al. (2021)), which strongly influence the Sun, the stars, the Universe?! This means that, unlike the holographic mechanism of quantum gravity on the Sun, where the magnetic flux tubes in strong fields ($\sim 4.1\cdot 10^{7}~G$ in the tachocline are identical to the magnetic core of the Sun), such physics of formation, emergence and “decay” of sunspots, which is considered an integral part of the convective dynamo (see e.g. Y. Fan & F. Fang (2014) and Fig. 21 in H. Hotta & H. Iijima (2020)), and, as a consequence, the formation of active regions from the uppermost convection zone to the corona (see e.g. F. Chen et al. (2017, 2022)), is still an unsolved problem of the 70-year evolution and the current state of solar and stellar dynamo theory (see e.g. P. Charbonneau & D. Sokoloff (2023) and P. Charbonneau (2020)), which, although good, cannot be on the Sun, and certainly not in our Universe!
@@
Finally, let us show that if the rise velocity of the buoyant magnetic tube, determined by the magnetic field values maximum ($B_{max}\sim 3.6\cdot 10^{7}~G$) and minimum ($B_{max}\sim 3.6\cdot 10^{6}~G$) of the magnetic field and the width of the “ring” $(a_{conv})_{max}\sim 3.7\cdot 10^{-4}~H_{p}$ or $(a_{conv})_{min}\sim 3.7\cdot 10^{-3}~H_{p}$, causes the MFT to appear on the surface in the form of sunspots, then the parameters of the universal vanBFF model will be associated with the magnetic cycles that are directly related to the observed features of the cycle of both double sunspot maxima (see Sect. 2.2.2) and the tilt angle of Joy’s law (see Sect. 2.2.1).
2.2 Physics of primary and secondary reconnection of magnetic flux tubes in the lower layers of the tachocline
@@
2.2.1 On the ascending magnetic tube and the formation of sunspots both near the tacho-cline with $\sim 10^{7}~G$ in the absence of Joy’s law and in the middle region with $\sim 10^{5}~G$ in the presence of Joy’s law
@@
The most interesting thing is that, in addition to the sunspots after the rise of the magnetic tube with $B\sim 10^{7}~G$ at all latitudes (see Fig. 9a and Fig. 10a) and the sunspots after the rise of the magnetic tube with $B\leqslant 2\cdot 10^{5}~G$ only at low latitudes (see Fig. 9b and Fig. 10c), which are the source of the double maxima of the 11-year solar cycles (see the full experimental evidence in Section 2.2.2), we are interested in the problem or, more precisely, the physics of the tilt angle of Joy’s law (V. G. Ivanov, 2012; M. Dasi-Espuig et al., 2010) and how it is related to the magnetic $\nabla\rho$-pumping?
@@
2.2.2 New physics of observed double maxima of sunspot cycles and solutions of time interval or Gnevyshev gap
@@
  • Secondly, it is an important holographic principle of quantum gravity. The most surprising fact is that the magnetic flux tube model at strong fields of the order of $10^{7}~G$ is a consequence of the thermomagnetic EN effect in the tachocline (see Fig. C.1a in V. D. Rusov et al. (2021)), which is the source of the so-called solar holographic mechanism of quantum gravity (see G. ’t Hooft (1993); L. Susskind (1995); J. Maldacena (2005); S. W. Hawking (2015b, a); S. W. Hawking et al. (2016). At the same time, we know that the hologram of the Sun, which is encoded in the magnetic field on its two-dimensional tachocline surface and is equivalent to the magnetic field (from the core to the tachocline) in the internal three-dimensional space of the Sun’s gravity, is in our case a remarkable source of the process with a continuous transformation of the toroidal field into a poloidal one (transformation $T\rightarrow P$), but not vice versa ($P\rightarrow T$) (see Fig. 2b and Eq. (2); also Fig. C.1a in V. D. Rusov et al. (2021)).

### v2

@@ I Introduction
On the other hand, we understand that the existence of the Parker-Biermann cooling effect is associated with the so-called thermomagnetic Ettingshausen-Nernst effect (see Apendix A in V. D. Rusov et al. (2021)). Due to the large temperature gradient in the tachocline, the thermomagnetic EN effect (A. V. Ettingshausen & W. Nernst, 1886; E. H. Sondheimer, 1948; L. J. Spitzer, 1956; Y. B. Kim & M. J. Stephen, 1969) creates electric currents that are inversely proportional to the strong magnetic field of the tachocline. As we showed earlier (V. D. Rusov et al., 2021), the toroidal magnetic field of tachocline by means of the thermomagnetic Ettingshausen-Nernst effect
@@
can indirectly prove that toroidal magnetic field of the tachocline is exactly equal to
$B_{tacho}^{Sun}=4.1\cdot 10^{7}~G,$ (2)
Using the strong MFT field, we are interested in the existence of dark matter axions identical to solar axions, which strongly affects the the conversion of axion origin photons in magnetic O-loops near the tachocline, and are thus related to the so-called thermomagnetic Ettingshausen-Nernst effect through the Parker-Biermann cooling effect (see Apendix A in V. D. Rusov et al. (2021)).
@@
Figure 1: An illustration of the main possible processes of a magnetically active star of the Sun type. (a) $\alpha$-effect, $\Omega$-effect and BL mechanism as components of the solar dynamo model. The $\Omega$-effect (blue) depicts the transformation of the primary poloidal field into a toroidal field by differential rotation. Regeneration of the poloidal field is then performed either by the $\alpha$-effect (top) or by the BL mechanism (yellow in the middle). In case of $\alpha$-effect, the toroidal field at the base of the convection zone is subject to cyclonic turbulence. In the BL mechanism, the main process of regeneration of the poloidal field (based on the $\Omega$-effect (blue)) is the formation of sunspots on the surface of the Sun from the rise of buoyant toroidal flux tubes from the base of the convection zone. The magnetic fields of these sunspots closest to the equator in each hemisphere diffuse and join, and the field due to the spots closer to the poles has a polarity opposite to the current that initiates rotation of the polarity. The newly formed polar magnetic flux is transported by the meridional flow to deeper layers of the convection zone, thereby creating a new large-scale poloidal field. Derived from S. Sanchez et al. (2014). (b) 11-year variations of sunspot magnetic fields as a component of the solar anti-dynamo. The formation of the thermomagnetic EN effect (see L. J. Spitzer (1962, 2006); V. Rusov et al. (2015)), emphasizes that this process is associated with the continuous transformation of toroidal field into poloidal one ($T\rightarrow P$ transformation), but not vice versa ($P\rightarrow T$). For this purpose, we note (seeV. D. Rusov et al. (2021)) that the modulations of the ADM halo in the Galactic Center (GC), which is an indicator of the periods of S-stars, e.g. S-102 with a period of about 11 years, lead (via vertical density waves from the black hole to the solar neighborhood) to modulations of the ADM density, gravitationally trapped in the interior of the Sun. This means that anticorrelation modulations of the ADM density control the solar luminosity or sunspot cycles, in which the sunspot variations are identical to the 11-year variations of the magnetic flux tubes (see Section II): maximum (red lines) and minimum (green-red lines).
@@
Taking into account the above, we structure our paper as follows. In Section II, we discuss the physics of nearly empty flux tubes and the connection with dark matter axions out of the solar core. In Section II.1, we present convective heating and magnetic buoyancy of flux tubes by means of axion origin photons. In Section II.2, we discuss the physics of primary and secondary reconnection of flux tubes in the lower layers and the features of the Joy’s law tilt angle. In Section II.2.2, we discuss the physics of the observed double maxima of sunspot cycles, where the first maximum does not depend on the tilt angle of the Joy’s law, while the second one definitely depends on the tilt angle of the Joy’s law. We also showed the relationship between the ADM density in the solar interior (or, for example, the number of anticorrelated sunspots) and the Gnevyshev gap. Finally, in Section III, we provide a summary and outlook for this paper.
II Physics of the theory of practically empty magnetic flux tubes at strong fields, $\sim 10^{7}~G$, at which there is a free path for photons of axion origin from the tachocline to the photosphere.
@@
In order for our alternative theory (for example, magnetic flux tubes in strong fields using photons of axion origin) to solve all its problems (see Sections II.1-II.2), it is necessary to solve the most important problem: In what way come sunspots, generated by a strong magnetic field in a flux tube from the base of the convection zone to the surface of the Sun, have 11-year variations without any dynamo?
The answer is very simple! Almost all physicists, and in particular dynamo physicists, know that in the Universe, and, of course, in our galaxy, there are ADM, and the highest density of ADM is located exactly at the center of the Galaxy, i.e. near the black hole. At the same time, we know that between the black hole and the disk there are also so-called amazing S-stars [^1] Let us remind from where the fairly young S-stars (Fig. 8a in (V. D. Rusov et al., 2021)) appear in our galaxy between the black hole and the disk. Since we know that in our galaxy, in addition to the supermassive black hole (SMBH), there is the so-called intermediate black hole (IMBH; see evidence (S. Takekawa et al., 2019)). If we take into account the evolution of binary SMBH-IMBH, then when one body (SMBH) exchanges partners of binary stars orbiting the IMBH (Fig. 8b in V. D. Rusov et al. (2021)), and through the extreme gravitational-tidal field, one star is captured (SMBH) and loses energy, while the other escapes, receives all this energy and, due to the hypervelocity of B-type stars – the IMBH sling (Fig. 8b in V. D. Rusov et al. (2021)), simply flies out of the galaxy. From here it becomes clear how in other galaxies a young S-star can end up between a black hole and a disk., which have a little more than 30 stars. The most interesting thing is that the closest S-star to the black hole is called S-102, which has an 11-yer period of around the black hole (see Fig. 6 and Fig. 8 in V. D. Rusov et al. (2021)). From here, we understand that when the period of S-102 is very close to the black hole, then the high speed of this star is such that the dark matter is practically not captured by this star. But when the elliptical star S-102 rotates farther from the black hole, then the star gravitationally captures these ADMs.
@@
Figure 3: Topological effects of magnetic reconnection inside the magnetic tubes with the “magnetic steps” (Fig. B.1 in V. D. Rusov et al. (2021)). The left panel shows the temperature and pressure change along the radius of the Sun from the tachocline to the photosphere (J. Bahcall & M. Pinsonneault, 1992), $L_{MS}$ is the height of the magnetic shear steps. At $R\sim 0.72~R_{Sun}$ the vertical magnetic field, which is developed from the horizontal part of the magnetic field (step) with the participation of the O-loop through the well-known Kolmogorov turbulent cascade (see Fig. 4), reaches $B_{z}\sim 3600$ T, and the magnetic pressure $p_{ext}=B^{2}/8\pi\simeq 5.21\cdot 10^{13}~erg/cm^{3}$ (J. Bahcall & M. Pinsonneault, 1992). The very cool regions along the entire convective zone caused by the Parker-Biermann cooling effect have the virtually zero internal gas pressure, i.e. the maximum magnetic pressure in the magnetic tubes. The narrow “purple” rings between the O-loop and the tube walls ($\rho_{ext}=\rho_{int}^{0}\gg\rho_{int}$) with the Parker-Bierman cooling effect inside, are a very important result of the existence of convective heating $(dQ/dt)_{2}$ in Section II.1.
On the other hand, high-energy photon fluxes coming from the radiation zone through the thin “ring” part of the magnetic tube, i.e. between the tachocline and the “ring” (or more precisely between the magnetic wall and the O-loop of the tube; see Figs. 3 and 4), and which is applied without the Parker-Biermann cooling effect, allows to determine the convective heating $(dQ/dt)_{2}$ (see Sect. II.1).
@@
One of the final goals of this section is to determine the general regularities of the theory of magnetic flux tubes in strong fields of the order of $\sim 10^{7}~G$ using photons of axion origin – the so-called universal Ballegooijen-Fan-Fisher model (Sec. II.1), which are generated by the magnetic buoyancy of practically empty tubes, and as a consequence, their rise from the tachocline to the solar surface (Figs. 3-5). Another goal is the essence of the physics of the primary and secondary reconnection of magnetic flux tubes in the lower layers of the tachocline (Sect. II.2), which is directly related to the observed features of the cycle of both double sunspot maxima (see Sect. II.2.2) and the tilt angle of Joy’s law (see Sect. II.2.1). Moreover, both effects (see Sect. II.1-II.2) are caused by the existence (of dark matter?) – solar axions generated in the core of the Sun.
II.1 Radiative $(dQ/dt)_{1}$ sec-reconnectionand convective $(dQ/dt)_{2}$ heatings and the buoyant rise of magnetic flux tubes and the phenomenon of solar axions
@@
at which we have a very important (see section II.2.1) value of
@@
at which we have a very important (see Sections II.1-II.2) value of
@@
Using simple calculations of equations (3) and (II.1) for MFTs, it is easy to show that with the help of the
@@
Ultimately, we call these Eqs. (II.1)-(18), in honor of these remarkable “solar” physicists, the universal van Ballegooijen-Fan-Fisher model (vanBFF model).
@@
where we used the MFT heating rate which consists of the radiative heating rate $(dQ/dt)_{1}$ and convective heating rate $(dQ/dt)_{2}$ (see Eqs. (II.1)–(19), Eqs. (13) and (19) in (Y. Fan & G. Fisher, 1996) and Eqs. (8)–(9) in (M. A. Weber & Y. Fan, 2015)); the approximate ratio $|\delta T|/T_{ext}\sim 1/\beta$; the parameter $\alpha_{1}^{2}\approx 5.76$ (Y. Fan & G. Fisher, 1996; M. A. Weber & Y. Fan, 2015); $F_{rad}$ is the radiative energy flux (see Eq. (15) in (Y. Fan & G. Fisher, 1996), and also Eq. (6) and Fig. 1 in (M. A. Weber & M. K. Browning, 2016)); $\nabla_{e}\sim 1.287\nabla_{rad}$ (see Table 2 in (H. C. Spruit, 1977)); $H_{p}/a$ is the factor for the lower convection zone (see (R. M. Roettenbacher et al., 2016)), where it was previously believed that $a=(\Phi/\pi B_{tacho})^{1/2}\leqslant 0.1H_{p}$ (see 3.2.1 in (R. M. Roettenbacher et al., 2016)) with an average value of typical magnetic flux of $\Phi\sim 10^{20}-10^{22}~Mx$ (Y. Fan et al., 1993).
@@
Recalling the words of Yuhong Fan that “it remains to be more thoroughly investigated” to what extent is strong toroidal magnetic field, control by means of the self-generating dynamo in the overshoot tachocline region, one must also remember the existence of the amazing 11-year variation of dark matter density (see e.g. V. D. Rusov et al. (2021)), and, for example, the holographic mechanism of quantum gravity (see S. W. Hawking (2015a, b); S. W. Hawking et al. (2016); V. D. Rusov et al. (2021)), which strongly influence the Sun, the stars, the Universe?! This means that, unlike the holographic mechanism of quantum gravity on the Sun, where the magnetic flux tubes in strong fields ($\sim 4.1\cdot 10^{7}~G$ in the tachocline are identical to the magnetic core of the Sun), such physics of formation, emergence and “decay” of sunspots, which is considered an integral part of the convective dynamo (see e.g. Y. Fan & F. Fang (2014) and Fig. 21 in H. Hotta & H. Iijima (2020)), and, as a consequence, the formation of active regions from the uppermost convection zone to the corona (see e.g. F. Chen et al. (2017, 2022)), is still an unsolved problem of the 70-year evolution and the current state of solar and stellar dynamo theory (see e.g. P. Charbonneau & D. Sokoloff (2023) and P. Charbonneau (2020)), which, although good, cannot be on the Sun, and certainly not in our Universe!
@@
Finally, let us show that if the rise velocity of the buoyant magnetic tube, determined by the magnetic field values maximum ($B_{max}\sim 3.6\cdot 10^{7}~G$) and minimum ($B_{max}\sim 3.6\cdot 10^{6}~G$) of the magnetic field and the width of the “ring” $(a_{conv})_{max}\sim 3.7\cdot 10^{-4}~H_{p}$ or $(a_{conv})_{min}\sim 3.7\cdot 10^{-3}~H_{p}$, causes the MFT to appear on the surface in the form of sunspots, then the parameters of the universal vanBFF model will be associated with the magnetic cycles that are directly related to the observed features of the cycle of both double sunspot maxima (see Sect. II.2.2) and the tilt angle of Joy’s law (see Sect. II.2.1).
II.2 Physics of primary and secondary reconnection of magnetic flux tubes in the lower layers of the tachocline
@@
II.2.1 On the ascending magnetic tube and the formation of sunspots both near the tacho-cline with $\sim 10^{7}~G$ in the absence of Joy’s law and in the middle region with $\sim 10^{5}~G$ in the presence of Joy’s law
@@
The most interesting thing is that, in addition to the sunspots after the rise of the magnetic tube with $B\sim 10^{7}~G$ at all latitudes (see Fig. 9a and Fig. 10a) and the sunspots after the rise of the magnetic tube with $B\leqslant 2\cdot 10^{5}~G$ only at low latitudes (see Fig. 9b and Fig. 10c), which are the source of the double maxima of the 11-year solar cycles (see the full experimental evidence in Section II.2.2), we are interested in the problem or, more precisely, the physics of the tilt angle of Joy’s law (V. G. Ivanov, 2012; M. Dasi-Espuig et al., 2010) and how it is related to the magnetic $\nabla\rho$-pumping?
@@
II.2.2 New physics of observed double maxima of sunspot cycles and solutions of time interval or Gnevyshev gap
@@
  • Secondly, the most surprising fact is that the magnetic flux tube model at strong fields of the order of $10^{7}~G$ is a consequence of the thermomagnetic EN effect in the tachocline (see Fig. C.1a in V. D. Rusov et al. (2021)), which is in our case a remarkable source of the process with a continuous transformation of the toroidal field into a poloidal one (transformation $T\rightarrow P$), but not vice versa ($P\rightarrow T$) (see Fig. 1b and Eq. (2); also Fig. C.1a in V. D. Rusov et al. (2021)).